Problem is
4/9 of 2 yards = 16
how do they get it

Answer by Alan3354(69443)

Problem is
4/9 of 2 yards = 16
--------------
That's not a problem, it's a statement.
No units are given for the 16, tho.
-----------
4/9 of 2 yards = 16 inches ?
4/9 of 72 inches = 16 inches ?
It doesn't.
4/9 of 72 inches is 32 inches.
---------------
English units are almost useless.